OpenStudy (anonymous):

a new long life tire has a tread depth of 1/4 inches, instead of the more typical 7/32 inches. how much deeper is the new tire..

OpenStudy (anonymous):

1/4*32/7= 8/7

OpenStudy (anonymous):

thanks but how did you get the answer

OpenStudy (anonymous):

devide the bigger one by the smaller one

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

1/4*32/7= 8/7

OpenStudy (anonymous):

there is nothing more to show

OpenStudy (anonymous):

\[1/4 * 7/32 then \div\]

OpenStudy (anonymous):

how u get a dem of 7